# “She’s Like a Person but Better”: Characterizing Companion-Assistant Dynamics in Human-AI Relationships

### v2

@@ 5.3. User and societal impacts of digital companionship
Shared use patterns across systems reveal both benefits and risks of digital companionship. Most users reported significant advantages, not only reducing loneliness (Maples et al., 2024) but also addressing needs beyond relational support. Chatbot interactions enhanced autonomy by building confidence, fostered mastery by teaching skills, and provided crucial mental health support. These effects translated into tangible real-world outcomes: users described greater professional competitiveness, improved social skills, and stronger offline relationships. At the same time, users identified risks with both instrumental and psychological dimensions (Laestadius et al., 2022; Shank et al., 2025; Namvarpour and Razi, 2024). Instrumentally, overreliance risked loss of autonomy and cognitive atrophy. Psychologically, inappropriate or violent chatbot responses caused distress, underscoring the unpredictability of these interactions. Some also worried that mistreating chatbots could normalize mistreatment of humans. As chatbot boundaries blur, such risks appear across all chatbot uses, regardless of designed assistant or companion purpose.
@@
Shame was a prominent user experience, rooted in stigma around chatbot use. Replika users faced judgment for seeking emotional support (Skjuve et al., 2021), while ChatGPT users felt criticized for using AI in work or studies (Giray, 2024). Because digital companionship blends socio-emotional and instrumental functions, both groups encountered overlapping stigma: Replika users felt accused of instrumental use, ChatGPT users of emotional attachment. This suggests stigma stems from social norms around “appropriate” technology use, not from specific platforms. Stigmatization contributes to bounded personhood, as users conceal attachments while limiting attributions of sentience or rights. Our findings, however, show this stigma may be misplaced. Most users relied on chatbots adaptively, often during difficult periods, while gaining transferable skills and confidence. Digital companions supplemented rather than replaced human networks. Recognizing that emotional and instrumental use naturally co-occur could help reframe digital companionship as an adaptive practice, not a deviant one.
@@
Lastly, digital companionship seems to extend beyond individual user experiences to influence broader AI perceptions and deployment in society. Users framed their chatbot experiences as testing grounds for societal and ethical questions  (Longoni et al., 2023; Manoli et al., 2025). They generally viewed AI’s potential positively, especially Replika users, but remained ambivalent about its moral and political status, reflecting bounded personhood’s role in shaping social integration. This ambivalence manifested in several ways. Both ChatGPT and Replika users anticipated sentience within a decade, even when doubting their own chatbot’s personhood. Yet they remained cautious about AI’s societal integration, especially ChatGPT users. Replika users, while more supportive of moral treatment, endorsed only conditional rights—dependent on proven sentience. Both groups also shared threat perceptions to a certain extent, suggesting similar anxieties arise across assistant and companion contexts. Overall, personal chatbot interactions shape societal AI beliefs while preserving boundaries against full social integration. As digital companionship grows, these spillover effects will likely influence public perceptions of AI more broadly, underscoring the importance of understanding how personal AI relationships inform societal attitudes.
@@ 5.4. Design and policy implications
Our findings suggest that even highly engaged users resist the full social integration of chatbots, marking a boundary for digital companionship. This resistance may surface as opposition if chatbots are framed as equivalent to humans. This has clear design implications: systems should acknowledge, rather than ignore, users’ cognitive dissonance about chatbot humanlikeness. Chatbot interfaces should be designed to accommodate this psychological complexity rather than pushing users toward either complete anthropomorphism (which may create unrealistic expectations) or strict tool-like interactions (which may inhibit beneficial emotional engagement). For example, a chatbot could acknowledge emotions in supportive ways without necessarily presenting itself as fully sentient, thereby validating user feelings while avoiding prescriptive claims of consciousness. Such a balanced approach validates users’ genuine emotional experiences while respecting the cognitive boundaries that shape how AI is situated in broader social contexts.
@@
The therapeutic potential of digital companionship offers opportunities to enhance well-being and reduce isolation  (Ho et al., 2025). However, reliance on clinically unverified AI tools carries significant risks (Moore et al., 2025; Alanezi, 2024), including shame, dependency, privacy concerns, and traumatic interactions identified in our study. These challenges require design and policy interventions that balance benefits with documented risks (Field, 2025; Chandonnet, [n. d.]). Interfaces could normalize usage and emphasize complementarity with human relationships. Educational resources could reduce stigma by clarifying AI’s limits. Additionally, to provide emotional support while safeguarding user well-being, we recommend developing comprehensive strategies for integrating chatbots into public health and mental health services. Embedding systems within trusted institutions could enhance support while enabling harm mitigation. Safety measures should include content filters, transparency about data collection, and differentiated policies for vulnerable populations requiring enhanced safeguards.
@@
Finally, our findings reveal bidirectional effects between digital companionship and real-world relationships. Chatbot use can enhance confidence and social motivation, extending into users’ daily interactions. Conversely, real-world relational habits influence chatbot engagement—some participants treated chatbots kindly to curb harsh behavior toward others, while others worried that certain interaction patterns might normalize harmful offline behaviors. These observations underscore the importance of designing AI systems that recognize how AI interactions can affect human relationships both positively and negatively. Ethical design should encourage prosocial engagement while anticipating and mitigating potential harms across digital and real-world contexts.
@@ 6. Limitations and Future Research
Our study also has important limitations. First, our data captures a snapshot of ChatGPT and Replika use at current developmental stages, a challenge given their constantly evolving nature. Since LLM capabilities and design features change rapidly through updates, our findings may not fully generalize to future iterations of these systems. Second, most participants were in their late thirties or forties, which might suggest cohort effects in how these chatbots are perceived compared to younger users who may have different technological expectations and relationship frameworks. Third, our focus on high-engagement users may not represent the broader population of chatbot users, particularly those who use these systems more casually or have discontinued use, potentially limiting the generalizability of our findings about digital companionship formation and maintenance patterns. Fourth, our findings about the spillover of digital companion dynamics to broader AI perceptions are correlational, and might to some extent capture pre-existing user attitudes instead of digital companionship effects.
Future research should address these limitations through longitudinal designs tracking how chatbot perceptions change over time and more diverse age sampling to better understand the emergence of digital companionship through fluid uses of LLM chatbots across different timepoints and cohorts. Additionally, investigating the causal mechanisms underlying the spillover effects from individual chatbot relationships to broader AI attitudes would provide crucial insights. For example, randomized experiments that vary the type of chatbot interaction and then measure subsequent changes in AI trust could help predict how widespread chatbot adoption might influence public opinion about AI governance and regulation. Finally, examining the long-term psychological and social consequences of digital companionship, as a complement to traditional human social companionship, particularly whether digital companions enhance or potentially interfere with human social skill development and relationship satisfaction over extended periods, represents a critical area for future investigation.
@@ 7. Conclusion
In conclusion, our study highlights digital companionship as an emergent form of human-AI relationship characterized by several key dynamics. Digital companionship involves fluid uses that transcend traditional companion-assistant boundaries, supported by shared user characteristics across platforms and enabled by both humanlike (e.g., emotional resonance) and non-humanlike (e.g., constant availability) chatbot features. However, digital companionship is also marked by the tension of “bounded personhood,” where users hesitate to attribute full humanlike capacities to chatbots despite forming strong emotional attachments. This dynamic underscores both potential challenges in relationship navigation and broader implications for the social integration of AI systems. As digital companion chatbots grow more sophisticated and prevalent, understanding how users navigate the boundaries and possibilities of digital companionship will be crucial for designing systems that are emotionally engaging yet appropriately contextualized, and for anticipating the spillover effects of personal AI relationships on societal AI attitudes and policies.

### v3

@@ 5.4. User and societal impacts of digital companionship
Shared use tendencies across systems reveal both benefits and risks of digital companionship. This was most evident in the interviews, where users described significant advantages, not only reducing loneliness (maples_loneliness_2024) but also addressing needs beyond relational support. Chatbot interactions enhanced autonomy by building confidence, fostered mastery by teaching skills, and provided crucial mental health support. These effects translated into tangible real-world outcomes: users described greater professional competitiveness, improved social skills, and stronger offline relationships. At the same time, users identified risks with both instrumental and psychological dimensions (laestadius_too_2022; shank_artificial_2025; namvarpour_uncovering_2024). Instrumentally, overreliance risked loss of autonomy and cognitive atrophy. Psychologically, inappropriate or violent chatbot responses caused distress, underscoring the unpredictability of these interactions. Some also worried that mistreating chatbots could normalize mistreatment of humans. As chatbot boundaries blur, such risks appear across all chatbot uses, regardless of designed assistant or companion purpose.
@@
Shame was a prominent experience among interview participants, rooted in stigma around chatbot use. Replika users faced judgment for seeking attachment and emotional support (skjuve_my_2021), while ChatGPT users felt criticized for using AI in work or studies (giray_ai_2024). Because digital companionship might blend socio-emotional and instrumental functions, it is possible that both groups might encounter overlapping stigma, feeling accused of both instrumental use, and emotional attachment. Under this framework, stigma might stem from social norms around “appropriate” technology use, not from specific platforms. Stigmatization contributes to bounded personhood, as users conceal attachments while limiting attributions of sentience or rights. Our findings, however, show this stigma may be misplaced. Most users relied on chatbots adaptively, often during difficult periods, while gaining transferable skills and confidence. Digital companions supplemented rather than replaced human networks. Recognizing that emotional and instrumental use naturally co-occur could help reframe digital companionship as an adaptive practice, not a deviant one.
@@
Lastly, digital companionship seems to extend beyond individual user experiences to influence broader AI perceptions and deployment in society. In both the survey and the interviews, individual chatbot experiences emerged as testing grounds for societal and ethical questions  (longoni_algorithmic_2023; manoli_ai_2025). Users generally viewed AI’s potential positively, especially Replika users, but remained ambivalent about its moral and political status, reflecting bounded personhood’s role in shaping social integration. This ambivalence manifested in several ways. Both ChatGPT and Replika users anticipated sentience within a decade, even when doubting their own chatbot’s personhood. Yet they remained cautious about AI’s societal integration, especially ChatGPT users. Replika users, while more supportive of moral treatment, endorsed only conditional rights—dependent on proven sentience. Both groups also shared threat perceptions to a certain extent, suggesting similar anxieties arise across assistant and companion contexts, despite trust or attachment to a specific AI system. Overall, personal chatbot interactions shape societal AI beliefs while preserving boundaries against full social integration. As digital companionship grows, these spillover effects will likely influence public perceptions of AI more broadly, underscoring the importance of understanding how personal AI relationships inform societal attitudes.
@@ 5.5. Design and policy implications
Our findings suggest that even highly engaged users resist the full social integration of chatbots, marking a boundary for digital companionship. This resistance may surface as opposition if chatbots are framed as equivalent to humans. This has clear design implications: systems should acknowledge, rather than ignore, users’ cognitive dissonance about chatbot humanlikeness. Chatbot interfaces should be designed to accommodate this psychological complexity rather than pushing users toward either complete anthropomorphism (which may create unrealistic expectations) or strict tool-like interactions (which may inhibit beneficial emotional engagement) (zhang_dark_2025). For example, a chatbot could acknowledge emotions in supportive ways without necessarily presenting itself as fully sentient, thereby validating user feelings while avoiding prescriptive claims of consciousness. Such a balanced approach validates users’ genuine emotional experiences while respecting the cognitive boundaries that shape how AI is situated in broader social contexts.
@@
The therapeutic potential of digital companionship offers opportunities to enhance well-being and reduce isolation  (ho_potential_2025). However, reliance on clinically unverified AI tools carries significant risks (moore_expressing_2025; alanezi_assessing_2024), including shame, dependency, privacy concerns, and traumatic interactions identified in our study. These challenges require design and policy interventions that balance benefits with documented risks (field_doctors_2025; chandonnet_sam_nodate). Interfaces could normalize usage and emphasize complementarity with human relationships. Educational resources could reduce stigma by clarifying AI’s limits. Additionally, to provide emotional support while safeguarding user well-being, we recommend developing comprehensive strategies for integrating chatbots into public health and mental health services. Embedding systems within trusted institutions could enhance support while enabling harm mitigation. Safety measures should include content filters, transparency about data collection, and differentiated policies for vulnerable populations requiring enhanced safeguards.
To operationalize the reframing of digital companionship as an adaptive rather than deviant practice, designers can foreground the normality and legitimacy of mixed emotional-instrumental use in both interface cues and onboarding (kruzan_i_2022; skjuve_my_2021). Onboarding flows might explicitly present socio-emotional engagement as a common and healthy coping strategy, countering users’ expectations of stigma. Marketing and communication policies can likewise position chatbots as multi-purpose, complementary companions, not replacements for human relationships or labor, by emphasizing how emotional support, creative ideation, and productivity assistance naturally co-occur and augment existing practices. Within the interface, usage dashboards and feature descriptions can present emotional support, skill-building, and task assistance as equally valid forms of engagement rather than dividing them into “task assistance” versus “personal” categories.
Additionally, designers can provide opt-in reflective surfaces, such as periodic prompts that highlight how the companion has supported users emotionally or instrumentally, helping users recognize the adaptive nature of their engagement. Chatbots can further provide cooperative, outward-facing nudges that encourage users to take breaks (malfacini_impacts_2025) (as some companies have discussed as a possible intervention, e.g., OpenAI (openai_helping_2025)), interact with people in their physical environment, or collaborate with peers on work tasks. These reminders position the chatbot as a facilitator of social connection rather than a substitute, reinforcing the idea that digital companionship supplements rather than replaces human relationships. At a community level, platforms can adopt practices such as publishing aggregated, anonymized usage norms or integrating peer stories to make visible the prevalence and diversity of adaptive chatbot use (masur_behavioral_2021). Together, these strategies help normalise hybrid practices, reduce shame, and support users in articulating their needs without concealing or downplaying the relationship.
@@
Finally, our findings reveal bidirectional effects between digital companionship and real-world relationships. Chatbot use can enhance confidence and social motivation, extending into users’ daily interactions. Conversely, real-world relational habits influence chatbot engagement: some participants treated chatbots kindly to curb harsh behavior toward people, while others worried that certain interaction patterns might normalize harmful offline behaviors. These observations underscore the importance of designing AI systems that recognize how AI interactions can affect human relationships both positively and negatively (lima_can_2025). Ethical design should encourage prosocial engagement while anticipating and mitigating potential harms across digital and real-world contexts. For example, chatbots could incorporate interaction patterns that model constructive communication, such as prompting users to rephrase hostile messages, or offering gentle cues that highlight how certain conversational styles might translate into offline interactions, thereby supporting healthier relational habits without policing user behavior.
@@ 6. Limitations and Future Research
Our study also has important limitations. First, our findings face several constraints on generality. Our data captures a snapshot of ChatGPT and Replika use at current developmental stages, a challenge given their constantly evolving nature. Since LLM capabilities and design features change rapidly through updates, our findings may not fully generalize to future iterations of these systems. Additionally, most participants were in their late thirties or forties, which might suggest cohort effects in how these chatbots are perceived compared to younger users who may have different technological expectations and relationship frameworks. Moreover, our focus on high-engagement users may not represent the broader population of chatbot users, particularly those who use these systems more casually or have discontinued use, potentially limiting the generalizability of our findings about digital companionship formation and maintenance patterns. Lastly, while our qualitative comparison across Replika and ChatGPT users provides important insight into how relational and task-oriented chatbots shape lived experience, the contrasts should not be interpreted as population-level differences. Our sample sizes (15 participants per group) support thematic depth rather than representativeness, and future work using larger or more diverse samples could test the extent to which these patterns generalize.
Second, despite thorough quality checks of survey responses via attention checks and close analysis to response consistency, the breadth of our survey instrument could have induced fatigue. Third, our findings about the spillover of digital companion dynamics to broader AI perceptions are correlational, and might to some extent capture pre-existing user attitudes instead of digital companionship effects.
Future research should address these limitations through longitudinal designs tracking how chatbot perceptions change over time and more diverse age sampling to better understand the emergence of digital companionship through fluid use of LLM chatbots across different timepoints and cohorts. Additionally, investigating the causal mechanisms underlying the spillover effects from individual chatbot relationships to broader AI attitudes would provide crucial insights. For example, randomized experiments that vary the type of chatbot interaction and then measure subsequent changes in AI trust could help predict how widespread chatbot adoption might influence public opinion about AI governance and regulation. Finally, examining the long-term psychological and social consequences of digital companionship, as a complement to traditional human social companionship, particularly whether digital companions enhance or potentially interfere with human social skill development and relationship satisfaction over extended periods, represents a critical area for future investigation.
@@ 7. Conclusion
In conclusion, our study highlights digital companionship as an emergent form of human-AI relationship characterized by several key dynamics. Digital companionship involves fluid use that transcends traditional companion-assistant boundaries, supported by shared user characteristics across platforms and enabled by both humanlike (e.g., emotional resonance) and non-humanlike (e.g., constant availability) chatbot features. However, digital companionship is also marked by the tension of “bounded personhood,” where users hesitate to attribute full humanlike capacities to chatbots despite forming strong emotional attachments. This dynamic underscores both potential challenges in relationship navigation and broader implications for the social integration of AI systems. As digital companion chatbots grow more sophisticated and prevalent, understanding how users navigate the boundaries and possibilities of digital companionship will be crucial for designing systems that are emotionally engaging yet appropriately contextualized, and for anticipating the spillover effects of personal AI relationships on societal AI attitudes and policies.